The Second International Symposium on Differential Geometry

July 4–6, 2003, Saga University, Saga 840-8502, Japan

Supported by Grant-in-Aid for Scientific Research from the Ministry of Education, Culture, Science and Technology

Programme

July 4 , Friday

14:00–15:00 Takashi Shioya

(Mathematical Institute, Tohoku University） Gromov-Hausdorff convergence of Sobolev spaces

15:10–16:10 Atsushi Fujioka

(Graduate School of Economics, Hitotsubashi University) Surfaces parametrized by a generalized Chebyshev net

16:30–17:30 Shin Nayatani

(Graduate School of Mathematics, Nagoya University) Combinatorial harmonic maps and superrigidity

July 5, Saturday

9:20–10:20 Mitsuhiro Itoh

(Institute of Mathematics, University of Tsukuba) 4-dimensional Goldberg conjecture and Seiberg-Witten theory

10:30–11:30 Yukihiro Mashiko

（Department of Mathematics, Saga University） Excess comparison theorems on Alexandrov surfaces

11:40–12:40 Jin-ichi Itoh

(Department of Mathematics, Kumamoto University） On the length of cut locus for finitely many points

14:00–15:00 Naoya Ando

（Department of Mathematics, Kumamoto University） An isolated umbilical point on a surface

15:10–16:10 Shigeo Kawai

(Department of Mathematics, Saga University) TBA

16:30–17:30 Kazuo Akutagawa

(Department of Mathematics, Shizuoka University) Yamabe metrics on cylindrical manifolds

July 6, Sunday

9:20–10:20 Miyuki Koiso

(Department of Mathematics, Kyoto University of Education） Stability for geometric variational problems with constraint

10:30–11:30 Reiko Miyaoka

(Graduate School of Mathematics, Kyushu University） TBA